Could someone please show me how to solve this? A cellular phone company offers a contract for which the C, in dollars, of the t minutes of telephoning is given by C=0.25(t-300)+42.95. Where it is assumed that t is less than or equal to 300 minutes. What time will keep the costs between $84.95 and $114.45 the telephoning time must be between _ and _ minutes
Answer by stanbon(75887)

Could someone please show me how to solve this? A cellular phone company offers a contract for which the C, in dollars, of the t minutes of telephoning is given by C=0.25(t-300)+42.95. Where it is assumed that t is less than or equal to 300 minutes. What time will keep the costs between $84.95 and $114.45 the telephoning time must be between _ and _ minutes
---------------------------
t <= 300
84.95 <= 0.25(t-300)+42.95 <= 114.45
---
Subtract 42.95 along the line to get:
42.50 <= 0.25(t-300) <= 71.5
---
Divide thru by 0.25 to get:
170 <= t-300 <= 286
---
Add 300 to get:
470 <= t <= 586
---
But t <=300
==============
Hummmmmm????
====================
Cheers,
Stan H.